A student pushes a loaded sled whose mass is 240 kg for a distance of 2.3 m over the
frictionless surface of a frozen lake. He exerts a horizontal force equal to 130 N. If the sled starts from
rest, what is its final velocity ?

Grade:11

1 Answers

SRC Jayanth
7 years ago
force= 130 N
mass= 240 kg
acceleration= 13/24 ms-2
displacement= 2.3 m
u=0ms-1
by the equations of motion
v2 – u2 = 2as
v2 – 02 = 2(13/24) 2.3
v2  = 2.49166
v = 1.5785013990068766064066755315403 ms-1